Jennifer White recently formed Southwest Land Design to fill a need for designing, installing and maintaining natural, desert oriented residential landscaping. In January of this year, its first month of operations, the following transactions occurred.

1. White invests $200,000 of her own money into the company. 2. Purchased an office building for $150,000 in cash. The building is valued at $110,000, and the remainder of the value is assigned to the land.

3. Signed a three-year promissory note at the bank for $125,000.

4. Purchased office equipment at a cost of $50,000, paying $10,000 down and agreeing to pay the remainder in 10 days.

5. Paid wages and salaries of $13,000 for the first half of the month. Office employees are paid twice per month.

6. Paid the balance on the office equipment from transaction #4.

7. Sold $54,000 of advertising during the first month. Customers have until the 15th of the following month to pay their bills.

8. Paid wages and salaries of $15,000 for the second half of the month.

9. Recorded $3,500 in commissions earned by the sales people during the month. They will be paid on the fifth of the following month.

10. White draws $10,000 from the company as her salary for the month.

Complete the accounting equation worksheet for transactions 1-10.

Total the worksheet and verify that Assets = Liabilities + Equity.

AND

In proper order and form, prepare an income statement, statement of owner’s equity

and balance sheet.

Southwest Land Design

Transactions for January

ASSETS

LIABILITIES

EQUITY

Accounts

Office

Accounts

Commissions

Notes

#

Cash

Receivable

Land

Equipment

Building

=

Payable

Payable

Payable

+

Amount

Description

1.  

=

+

2.  

=

+

3.  

=

+

4.  

=

+

5.  

=

+

6.  

=

+

7.  

=

+

8.  

=

+

9.  

=

+

10.  

=

+

Totals

=

+
Answer #1
ASSETS LIABILITIES EQUITY
Accounts Office Accounts Commissions Notes
# Cash Receivable Land Equipment Building = Payable Payable Payable + Amount Description
1.   200000 = + 200000 Owner brings in capital in the form of cash
2.   -150000 40000 110000 = + Company buys land and buildings for cash
3.   125000 = 125000 + Borrows on signing promissory note
4.   -10000 50000 = 40000 + Purchases office equipment by paying cash for $10000 and on account for the balance.
5.   -13000 = + -13000 Pays wages for the first half of the month
6.   -40000 = -40000 + Settles accounts payable
7.   54000 = + 54000 Renders advertising services on account
8.   -15000 = + -15000 Pays wages for the second half of the month
9.   = -3500 + -3500 Employees earn commissions for the month; to be paid in the following month
10.   -10000 = + -10000 White withdraws cash as salary for the month
87000 54000 40000 50000 110000 0 -3500 125000 212500
Accounting equation: Assets = Liabilities+Equity
Assets = 87000+54000+40000+50000+110000 = 341000
Liabilities = 3500+125000+212500 = 341000
INCOME STATEMENT
Service revenue 54000
Expenses:
Wages 28000
Sales commission 3500
Salary to White 10000
Total expenses 41500
Net income 12500
STATEMENT OF OWNER'S EQUITY
Beginning balance 0
Cash brought in 200000
Net income 12500 212500
Ending balance 212500
BALANCE SHEET
ASSETS
Current Assets:
Cash 87000
Accounts receivable 54000
Total current assets 141000
Non-current assets:
Equipment 50000
Buildings 110000
Land 40000
Total non-current assets 200000
Total assets 341000
TOTAL LIABILITIES AND OWNER'S EQUITY
Current liabilities:
Accounts payable 0
Commissions payable 3500
Total current liabilities 3500
Long term notes payable 125000
Total liabilities 128500
Owner's equity 212500
Total liabilities and owner's equity 341000